# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v2

@@ 4.2.2 The Validity of Dysca
In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks. We calculate the Spearman’s rank correlation coefficient [63] $\rho$ and the Kendall rank correlation coefficient [64] $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench [39], OCRBench [52] and SeedBench-2 [24]. Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation. Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types. We then calculate the correlation of model evaluation rankings within this intersection. The results are shown in the first row of Tab. 5. For the MMbench [39] and OCRBench [52], both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6. However, the correlation for SeedBench-2 [24] is not as strong. Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only. As shown in the second row of Tab. 5, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs. 0.64 for $\rho$ and 0.43 vs. 0.62 for $\tau$). The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
To further explore the the impact of image styles on evaluation results, we present the average scores across all subtasks for each of the 51 styles in Fig. 5. We observe slight score differences across all styles. In the case of realistic styles such as "iPhone photo", all LVLMs perform better compared to other image styles. The LVLMs also exhibit better performance on unrealistic but common styles like "expressionist". However, for unrealistic and less common styles such as "gothic", all models show relatively poor performance. The results reveal that the gap between Dysca and non-synthesis benchmarks primarily stems from the more diverse range of image styles, making Dysca a more comprehensive benchmark for assessing the perception ability compared to previous benchmarks.
Figure 5: Illustration of each model’s performance across 51 image styles, where the darker colors represent higher scores. The representative styles are colored with non-black font. Realistic styles are shown in red font, unrealistic but common styles are displayed in yellow font, and unrealistic and less common styles are represented in blue font.
@@ 5 Conclusion
In this paper, we purpose Dysca, a dynamic and scalable benchmark for evaluating perception ability of Large Vision Language Models (LVLMs). Dysca consists of 617K Vision-language QA pairs, covering 20 perceptual subtasks, 4 image scenarios and 3 question types. We conduct the experiment on 8 advanced open-source LVLMs with 10 checkpoints, revealing the insightful weakness of current LVLMs when facing different question types, image styles and image conditions. Experiments demonstrate the validity on evaluating LVLMs by using synthesis images.
@@ Acknowledgement
This work is partially supported by National Key R&\&&D Program of China (No. 2021YFC3310100), Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680000), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS.

### v3

@@ 4.4 The Validity of Dysca
In this section, we investigate on the evaluation gap between Dysca and non-synthesis benchmarks. We calculate the Spearman’s rank correlation coefficient (Spearman, 1904) $\rho$ and the Kendall rank correlation coefficient (KENDALL, 1938) $\tau$ between the evaluation ranking of Dysca under clean scenario with the non-synthesis benchmark’s evaluation ranking, i.e., MMBench (Liu et al., 2023c), OCRBench (Liu et al., 2024) and SeedBench-2 (Li et al., 2023b). Both coefficient generate a score in the range of [-1,1], where 1 represents a perfect positive correlation, -1 represents a perfect negative correlation, and 0 represents no correlation. These coefficients are typical tools for measuring the correlation between variables in statistics. When the absolute value of either coefficient exceeds 0.6, it is considered to indicate a significant correlation  (Akoglu, 2018). Specifically, we intersect our Dysca with current benchmarks based on the perceptual subtasks, evaluation models and evaluation question types. We then calculate the correlation of model evaluation rankings within this intersection. The results are shown in the first row of Tab. 4. For the MMbench (Liu et al., 2023c) and OCRBench (Liu et al., 2024), both metrics show the high correlation, with $\rho$ and $\tau$ higher than 0.6. However, the correlation for SeedBench-2 (Li et al., 2023b) is not as strong. Considering that SeedBench-2 only contains realistic images, we conduct additional experiments using the evaluation ranks on our realistic style images only. As shown in the second row of Tab. 4, the correlation results of SeedBench-2 significantly improve (i.e., 0.46 vs. 0.64 for $\rho$ and 0.43 vs. 0.62 for $\tau$). The correlation with OCRBench also improves to 1, demonstrating the validity of using synthetic datasets for evaluation LVLMs.
Figure 7: The KMMD distance between each benchmarks, with darker colors indicating larger distances.
Besides, we calculate the data distribution distance between each benchmark to prove the low distance distribution between Dysca and non-synthesis benchmarks. We select CCBench (Liu et al., 2023c), COCO-Val  (Lin et al., 2014), MMVet (Yu et al., 2023), MMBench (Liu et al., 2023c), MME (Fu et al., 2023), MMStar (Chen et al., 2024), OCRBench (Liu et al., 2024) and ScienceQA (Lu et al., 2022). The reason why we choose these benchmark is that they have been widely used in evaluating LVLMs. We use Kernel Maximum Mean Discrepancy (KMMD) (Schölkopf et al., 2007) to measure the distribution distance. Specifically, we randomly sample 3,000 images from each benchmark (if the scale of the benchmark less than 3000, we use all that data) and utilize CLIP  (Radford et al., 2021) to encode these images. Then, we calculate the KMMD value using an RBF kernel between each pair. The results are shown in Fig. 7, with darker colors indicating larger distances. Each row represents the value of KMMD between two benchmarks. The value of last row denotes the average value of KMMD. As can be seen, the distribution distance between Dysca and real-image benchmarks ranks in the middle compared to all other benchmarks. This demonstrates that the distribution between our benchmark and real benchmarks is similar, indicating that the evaluation results can effectively reflect the model’s performance in real-world scenarios.
@@ 5 Conclusion
In this paper, we purpose Dysca, a dynamic and scalable benchmark for evaluating perception ability of Large Vision Language Models (LVLMs). Dysca consists of 617K Vision-language QA pairs, covering 20 perceptual subtasks, 4 image scenarios and 3 question types. We conduct the experiment on 24 advanced open-source LVLMs and 2 closed-source LVLMs, revealing the insightful weakness of current LVLMs when facing different question types, image styles and image conditions. Experiments demonstrate the validity on evaluating LVLMs by using synthesis images.
@@ Acknowledgement
This work is partially supported by Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0680202), Beijing Nova Program (20230484368), Suzhou Frontier Technology Research Project (No. SYG202325), and Youth Innovation Promotion Association CAS.